how many milliliters of 12.0 M HCl are needed to prepare 875.0 ml of 1.00 M HCl(aq)
Answer #1

use dilution formula

M1*V1 = M2*V2

1---> is for stock solution

2---> is for diluted solution

Given:

M1 = 12 M

M2 = 1 M

V2 = 875 mL

use:

M1*V1 = M2*V2

V1 = (M2 * V2) / M1

V1 = (1*875)/12

V1 = 72.9 mL

Answer: 72.9 mL